OpenStudy (anonymous):

what is the integral of sin^2x dx?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

sin^2(x)=(1-cos(2x))/2

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

you're welcome :). And if you integrate sin^2(x) over one or more full "loops" there is a trick since the average of sin^2(x) over a period is 1/2

OpenStudy (anonymous):

for example: \[\int\limits_{0}^{\pi}\sin^2(x)dx=\frac{\pi \int\limits_{0}^{\pi}\sin^2(x)dx}{\pi}=\frac{\pi}{2}\]

OpenStudy (anonymous):

I am an engineer:) I have to solve this basic equation:)

OpenStudy (anonymous):

allright then I geuss don't worry about it:)

OpenStudy (anonymous):

I have calculus final today and I forgot the integral of sin^2x:D

OpenStudy (anonymous):

oh allright well you can just use the trig identity. integral of cos^2 is similar: cos^2(x)=(1+cos(2x))/2

OpenStudy (anonymous):

these are just the trig identities you still have to integrate by the way

OpenStudy (anonymous):

int sin^2(x)= (x/2)-(sin(2x)/4)+C